rectangular room 20 feet long 18 feet wide whats the perimeter in meters
Answer by checkley77(12844)

20 ft=20/3=6.667 meters
18 ft=18/3=6 meters.
2*6.667+2*6=p
13.33+12=25.33 meters ans.